OpenStudy (lina17):

2|x-6|+14=38 step by step please

OpenStudy (anonymous):

ok step by step

OpenStudy (anonymous):

you have to get the absolute value by itself first first step therefore is to subtract \(14\) from both sides what do you get?

OpenStudy (lina17):

you get 2|x-6|=24

OpenStudy (anonymous):

yes you do next step in getting the absolute value by itself is to divide both sides by 2 now what is it?

OpenStudy (lina17):

you get x-6=12

OpenStudy (anonymous):

actually you should still have an absolute value there on the left

OpenStudy (lina17):

?

OpenStudy (anonymous):

you started with \[2|x-6|=24\] right?

OpenStudy (lina17):

yes

OpenStudy (anonymous):

if you divide by 2, the absolute value sign does not go away, it should still be there

OpenStudy (lina17):

|x-6|=12

OpenStudy (anonymous):

that's better

OpenStudy (anonymous):

because the second to last step is to get rid of the absolute value sign there are two numbers whose absolute value is 12 in is 12 the other is -12 so you have to solve two equalities and get two answers one is to solve \[x-6=12\] and the other is to solve \[x-6=-12\]

OpenStudy (lina17):

so would i divide

OpenStudy (anonymous):

no you would add 6 to both sides

OpenStudy (lina17):

oh ok

OpenStudy (lina17):

X-12=18

OpenStudy (lina17):

X-12= 16

OpenStudy (anonymous):

one answer is \(12+6=18\) yes

OpenStudy (lina17):

i mean x-12= -6

OpenStudy (anonymous):

the other is \(x=-6\) right

OpenStudy (lina17):

can you help me on one more problem

OpenStudy (anonymous):

sure

OpenStudy (lina17):

its |x|-8=-5

OpenStudy (lina17):

@satellite73